# Machine Translation Models are Zero-Shot Detectors of Translation Direction

### v2

@@ Abstract
Detecting the translation direction of parallel text has applications for machine translation training and evaluation, but also has forensic applications such as resolving plagiarism or forgery allegations. In this work, we explore an unsupervised approach to translation direction detection based on the simple hypothesis that , motivated by the well-known simplification effect in translationese or machine-translationese. In experiments with massively multilingual machine translation models across 20 translation directions, we confirm the effectiveness of the approach for high-resource language pairs, achieving document-level accuracies of 82–96% for NMT-produced translations, and 60–81% for human translations, depending on the model used.111Code and demo are available at https://github.com/ZurichNLP/translation-direction-detection
@@
Michelle Wastl Jannis Vamvas Rico Sennrich Department of Computational Linguistics, University of Zurich michelle.wastl@uzh.ch, {vamvas,sennrich}@cl.uzh.ch
@@ 1 Introduction
Previous work has addressed translation direction detection with feature-based approaches, using features such as n-gram frequency statistics and POS tags for classification Kurokawa et al. (2009); Volansky et al. (2013); Sominsky and Wintner (2019) or unsupervised clustering Nisioi (2015); Rabinovich and Wintner (2015). However, these methods require a substantial amount of text data, and cross-domain differences in the statistics used can overshadow differences between original and translationese text.
In this work, we explore the unsupervised detection of translation directions purely on the basis of a neural machine translation (NMT) system’s translation probabilities in both directions. As illustrated in Figure 1, we hypothesize that , which, if it generally holds, would allow us to infer the original translation direction.
@@
We test the approach on 20 translation directions, experimenting with 3 multilingual NMT models to predict the translation probabilities of human translations, NMT-produced translations, and pre-neural translations. We find that the approach detects the translation direction of human translations with an accuracy of 66% on average on the sentence level, and 80% for documents with 10 sentences. For the output of neural MT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
@@
We demonstrate that the approach is effective for detecting the original translation direction of neural translations, and to a lesser extent, human translations in a variety of high-resource language pairs.
@@ 2.1 Translation (Direction) Detection
In an ideal scenario where large-scale annotated in-domain data is available, high accuracy can be achieved in translation direction detection at phrase and sentence level by training supervised systems based on various features such as word frequency statistics and POS n-grams Sominsky and Wintner (2019).
To reduce reliance on in-domain supervision, unsupervised methods that rely on clustering and consequent cluster labelling have also been explored for the related task of translationese detection Rabinovich and Wintner (2015); Nisioi (2015). One could conceivably perform translation direction detection using similar methods, but this has the practical problem of requiring an expert for cluster labelling, and poor open-domain performance. In a multi-domain scenario, Rabinovich and Wintner (2015) observe that clustering based on features proposed by Volansky et al. (2013) results in clusters separated by domain rather than translation status. They address this by producing clusters, being the number of domains in their dataset, and labelling each. Clearly, labelling becomes more costly as the number of domains increases, which might limit applicability to an open-domain scenario.
In contrast, we hypothesize that comparing translation probabilities remains a valid strategy across domains, and requires no resources other than NMT models that are competent for the respective language pair and domain.
@@ 2.2 Translation Probabilities
So far, translation probabilities have been used in tasks such as noisy parallel corpus filtering Junczys-Dowmunt (2018), machine translation evaluation Thompson and Post (2020), and paraphrase identification Mallinson et al. (2017); Vamvas and Sennrich (2022). Those tasks all involve the comparison of parallel text and require a high level of attention to stylistic detail, for which the translation probabilities have proven useful in a highly data-efficient way by using existing machine translation models to generate the probabilities.
@@ 3.3 On Directional Bias
Ideally, we would want to match true value, i.e. equal 0.5 in a data set where gold directions are balanced. To allow for a cross-lingual comparison of bias despite varying data balance, we measure bias via the difference in accuracy between the two gold directions. An unbiased model should have similar accuracy in both directions. An extremely biased model that always predicts would achieve perfect accuracy on the gold direction , and zero accuracy on the reverse gold direction . We will report the bias as follows:
@@ 4 Experiments: Models and Data
We experiment with three massively multilingual machine translation models: M2M-100-418M (Fan et al., 2021), SMaLL-100 (Mohammadshahi et al., 2022), and NLLB-200-1.3B (NLLB Team et al., 2022).
The models are architecturally similar, all being based on the Transformer architecture Vaswani et al. (2017), but they differ in the training data used, number of languages covered, and model size, and consequently in translation quality. The comparison allows conclusions about how sensitive our method is to translation quality – NLLB-200-1.3B yields the highest quality of the three (Tiedemann and de Gibert, 2023), but we also highlight differences in data balance. English has traditionally been dominant in the amount of training data, and all three models aim to reduce this dominance in different ways, for example via large-scale back-translation Sennrich et al. (2016) in M2M-100-418M and NLLB-200-1.3B. SMaLL-100 is a distilled version of the M2M-100-12B model, and samples training data uniformly across language pairs.
We test the approach on datasets from the WMT news/general translation tasks from WMT16 Bojar et al. (2016), WMT22 Kocmi et al. (2022), and WMT23 Kocmi et al. (2023), which come annotated with document boundaries and the original language of each document. We also experiment with a part of the FLORES-101 dataset Goyal et al. (2022) to test the approach on indirect translations, where English was the original language of both sides of the parallel text. We divide the data into subsets based on several categorisations:
Type of translation: we distinguish between human translations (HT), which consist of (possibly multiple) reference translations, neural translation systems (NMT) (WMT 2016; 2022–2023), and phrase-based or rule-based pre-neural systems from WMT 2016 (pre-NMT) as a third category.
Directness: Given that the WMT data are direct translations from on side of the parallel text to the other, we perform additional experiments on translations for 4 FLORES language pairs (BengaliHindi, CzechUkrainian, GermanFrench, XhosaZulu). This allows us to analyze the behavior of our approach on indirect sentence pairs where both source and reference are translations from a source in a third language (in this case English).
We use HT and NMT translations from WMT16 as a validation set, and the remaining translations for testing our approach. Table 1 shows test set statistics for our main experiments.
@@
As an additional “real-world” dataset, we use 86 parallel sentences in German and English from a publicly documented plagiarism allegation case, in which translation-based plagiarism was the main focus (Ebbinghaus, 2022; Zenthöfer, 2022). The sentences were extracted from aligned excerpts of both the PhD thesis and the alleged source that were presented in a plagiarism analysis report (Weber, 2022). We extracted the sentences with OCR and manually checked for OCR errors. This dataset introduces an element of real-world complexity, as it involves translations that might not adhere strictly to professional standards nor does it have to fit into either of the aforementioned categories of translation strategy. It also provides a unique opportunity to test the robustness and adaptability of the translation direction detection system in a scenario that extends beyond controlled environments.
@@ 5.1 Sentence-level Classification
Table 2 compares the results for human translations across all models. As a general result, we find that it is not NLLB, but M2M-100 that on average yields the best results for human translations on the sentence level, with SMaLL-100 a very close second. Hence, we report results for M2M-100 in further experiments, and report performance of the other models in the Appendix.
A second result is that the translation detection works best for neural translations (75.0% macro-average), second-best for human translations (66.5% macro-average), and worst for pre-neural (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that the NMT systems we use tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems. In practice, one could pair our detection method with a monolingual model to identify such low-quality outputs.
@@ 5.2 Directional Bias
Analyzing Table 2 for directional bias, we observe that M2M-100 is especially biased in the directions defr () and zhen (). While we expected a general bias towards xen due to the dominance of English in training data, we find that the direction and strength of the bias vary across language pairs and models. An extreme result is NLLB for enzh, with towards zhen.
We leave it to future work to explore whether bias can be reduced via different normalization, a language pair specific bias correction term, or different model training. At present, our recommendation is to be mindful in the choice of NMT model and to perform validation before trusting the results of a previously untested NMT model for translation direction detection.
5.3 Indirect Translations
With an experiment on the English-original FLORES data, we evaluate our approach on the special case that neither side is the original. As shown in Table 5, our approach yields relatively balanced predictions on human translations for CzechUkrainian and XhosaZulu, predicting each direction a roughly equal number of times. For GermanFrench, we again find that the model predicts the defr direction much more frequently than the reverse direction, reflecting the high directional bias of the model for this language pair.
@@ 5.4 Qualitative Analysis
A qualitative comparison of sources and translations, as illustrated in Table 6, reveals that factors such as normalization, simplification, word order interference, and sentence length influence the detection of translation direction. In Example 1, an English HT translates the German source fully, while the NMT omits half of the content, showing a high degree of simplification. Our method recognizes the simplified NMT version as a translation but not the more literal HT. Example 2 demonstrates varying degrees of normalization: the first translation corrects typos, whereas the second shows stronger normalization with added capitalization and punctuation but also exhibits more interference due to closer adherence to the source’s lexical choice and copying the trailing quotation mark. Both translations are detected, but the second presents a higher probability difference. The third example indicates that translations exhibiting normalization, simplification, and interference to a higher degree are more likely to be identified. In Example 4, source language interference in terms of word order and choice significantly impacts the detection; the more literal translation mirroring the source’s word order is recognized, while the more liberal translation is not. Finally, Example 5 highlights challenges with short sentences: The German phrase Mit freundlichen Grüßen is fairly standardized, while its French equivalents can vary in use and context, adding to the ambiguity and affecting the probability distribution in NMT. Hence, our approach fails to identify any of the French translations without additional context.
Misclassified short sentences as in Example 5 are not a rarity in our experiments. Our findings show that reliable detection of translation direction, with an average accuracy exceeding 50%, is consistently attained for all language pairs we tested starting at a sentence length between 50 and 60 characters.555We used SMaLL-100 for this analysis. Additionally, we observed a trend where the accuracy of direction detection increases as the length of the sentences grows. This aligns with previous unsupervised approaches, which also documented higher accuracy the larger the text chunks that were used, although there, reliable results were reported on a more extreme scale starting from text chunks with a length of 250 tokens Rabinovich and Wintner (2015).
5.5 Document-Level Classification
Accuracy scores for document-level results by M2M-100 (best-performing system at sentence level) are presented in Tables 8 (NMT) and 7 (HT). We consider documents with at least 10 sentences, and language pairs with at least 100 such documents in both directions.

### v3

@@ Abstract
Detecting the translation direction of parallel text is useful not only for machine translation training and evaluation but also has forensic applications, such as resolving plagiarism or forgery allegations. In this work, we explore an unsupervised approach to translation direction detection based on the simple hypothesis that , motivated by the well-known simplification effect in translationese or machine-translationese. In experiments with multilingual machine translation models across 20 translation directions, we confirm the effectiveness of the approach for high-resource language pairs, achieving document-level accuracies of 82–96% for NMT-produced translations, and 60–81% for human translations, depending on the model used.111Code and demo are available at https://github.com/ZurichNLP/translation-direction-detection
@@
Michelle Wastl Jannis Vamvas Rico Sennrich Department of Computational Linguistics, University of Zurich {wastl,vamvas,sennrich}@cl.uzh.ch
@@ 1 Introduction
Previous work has addressed translation (direction) detection with feature-based approaches, using features such as n-gram frequency statistics and POS tags for classification Kurokawa et al. (2009); Volansky et al. (2013); Sominsky and Wintner (2019) or unsupervised clustering Nisioi (2015); Rabinovich and Wintner (2015). However, these methods require a substantial amount of text data, and cross-domain differences in the statistics used can overshadow differences between original and translationese text.
In this work, we explore the unsupervised detection of translation directions purely on the basis of a neural machine translation (NMT) system’s translation probabilities. As illustrated in Figure 1, we hypothesize that , which, if it generally holds, would allow us to infer the original translation direction.
@@
We test the approach on 20 translation directions, experimenting with 3 massively multilingual NMT models to predict the translation probabilities of human translations, NMT-produced translations, LLM-generated translations, and pre-neural machine translations. We find that the approach detects the translation direction of human translations with an accuracy of 66% on average on the sentence level, and 80% for documents with 10 sentences. For the output of neural MT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
To compare our unsupervised approach to a supervised baseline, we implement a modernized version of the approach proposed by Sominsky and Wintner (2019). A controlled comparison shows that a supervised approach can outperform our unsupervised one under ideal conditions (with in-domain labelled training data) for human translations, but performance drops in a cross-domain setting. Notably, for NMT-produced translations, the unsupervised approach remains competitive even when tested within the same domain.
@@
We demonstrate that the approach is effective for detecting the original translation direction of neural machine translations, and to a lesser extent, human translations in a variety of high-resource language pairs.
@@ 2.1 Translation (Direction) Detection
In an ideal scenario where large-scale annotated in-domain data is available, high accuracy can be achieved in translation direction detection at phrase and sentence level by training supervised systems based on various features such as word frequency statistics, POS n-grams or text embeddings Sominsky and Wintner (2019).
To reduce reliance on in-domain supervision, unsupervised methods that rely on clustering and consequent cluster labelling have also been explored for the related task of translationese detection Rabinovich and Wintner (2015); Nisioi (2015). One could conceivably perform translation direction detection using similar methods, but this has the practical problem of requiring an expert for cluster labelling, and poor open-domain performance. In a multi-domain scenario, Rabinovich and Wintner (2015) observe that clustering based on features proposed by Volansky et al. (2013) results in clusters separated by domain rather than translation status. They address this by producing clusters, being the number of domains in their dataset, and labelling each. Clearly, labelling becomes more costly as the number of domains increases, which limits applicability to an open-domain scenario.
In contrast, we hypothesize that comparing translation probabilities remains a valid strategy across domains, and requires no resources other than NMT models that are competent for the respective language pair.
@@ 2.2 Translation Probabilities
Previous work has leveraged translation probabilities for tasks such as noisy parallel corpus filtering Junczys-Dowmunt (2018), machine translation evaluation Thompson and Post (2020), and paraphrase identification Mallinson et al. (2017); Vamvas and Sennrich (2022). Those approaches analyze translation probabilities symmetrically in two directions, which is also the case in this work.
@@ 3.3 On Directional Bias
To allow for a cross-lingual comparison of bias despite varying data balance, we measure bias via the difference in accuracy between the two gold directions. An unbiased model should have similar accuracy in both. An extremely biased model that always predicts will achieve perfect accuracy on the gold direction , and zero accuracy on the reverse gold direction . We will report the bias as follows:
@@ 4 Experiments: Models and Data
4.1 Unsupervised Setting
We experiment with three multilingual machine translation models: M2M-100-418M (Fan et al., 2021), SMaLL-100 (Mohammadshahi et al., 2022), and NLLB-200-1.3B (NLLB Team et al., 2024).
The models are architecturally similar, all being based on the Transformer architecture Vaswani et al. (2017), but they differ in the training data used, number of languages covered, and model size, and consequently in translation quality. The comparison allows conclusions about how sensitive our method is to translation quality – NLLB-200-1.3B yields the highest translation quality of the three (Tiedemann and de Gibert, 2023), but we also highlight differences in data balance. English has traditionally been dominant in the amount of training data, and all three models aim to reduce this dominance in different ways, for example via large-scale back-translation Sennrich et al. (2016) in M2M-100-418M and NLLB-200-1.3B. SMaLL-100 is a distilled version of the M2M-100-12B model, and samples training data uniformly across language pairs.
@@ 4.1 Unsupervised Setting
We test the approach on datasets from the WMT news/general translation tasks from WMT16 Bojar et al. (2016), WMT22 Kocmi et al. (2022), and WMT23 Kocmi et al. (2023), which come annotated with document boundaries and the original language of each document. We include different years of WMT data not only to test the approach for different translation directions and translation types but also to rule out test set contamination for experiments with models, for which the test set predates the model release. The results for the individual WMT datasets are reported in Appendix F. We also experiment with a part of the FLORES-101 dataset Goyal et al. (2022) to test the approach on indirect translations, where English was the original language of both sides of the parallel text. We divide the data into subsets based on several categorizations:
Type of translation: we distinguish between human translations (HT), which consist of (possibly multiple) reference translations, neural translation systems (NMT) (WMT 2016; 2022–2023; (incl. translations from machine translation systems specifically and, to a lesser extent, translations produced by large language models (LLMs)), and phrase-based or rule-based pre-neural systems from WMT 2016 (pre-NMT) as a third category.
Directness: Given that the WMT data are direct translations from one side of the parallel text to the other, we perform additional experiments on translations for 6 FLORES language pairs (BengaliHindi, CzechUkrainian, GermanFrench, GermanHindi, ChineseFrench, and XhosaZulu). This allows us to analyze the behavior of our approach on indirect sentence pairs where both sequences are translations from a third language (in this case, English).
HT | NMT | ||||||
LP | Avg. | Avg. | |||||
encs | supervised | 64.70 | 71.36 | 68.03 | 65.43 | 71.98 | 68.71 |
unsupervised | 68.85 | 65.19 | 67.02 | 71.87 | 78.30 | 75.09 | |
enru | supervised | 56.65 | 81.74 | 69.19 | 61.72 | 82.97 | 72.35 |
unsupervised | 71.81 | 54.05 | 62.93 | 76.91 | 71.98 | 74.44 | |
ende | supervised | 87.05 | 55.29 | 71.17 | 89.52 | 61.08 | 75.30 |
unsupervised | 56.38 | 67.44 | 61.91 | 62.69 | 85.27 | 73.98 | |
Macro Avg. | supervised | 69.46 | 69.47 | 69.46 | 72.22 | 72.01 | 72.12 |
unsupervised | 65.68 | 62.23 | 63.95 | 70.49 | 78.51 | 74.50 |
We use HT and NMT translations from WMT16 as a validation set and the remaining translations for testing our approach. Table 1 shows test set statistics for our main experiments.
@@
To demonstrate our approach on “real-world” data, we use text pertaining to a publicly documented plagiarism allegation case, in which translation-based plagiarism was the main focus (Ebbinghaus, 2022; Zenthöfer, 2022). We use 86 parallel segments in German and English from aligned excerpts of both the PhD thesis and the alleged source that were presented in a plagiarism analysis report (Weber, 2022). We extracted the segments with OCR and manually checked for OCR errors. An example translation from this dataset is given in Appendix I.
4.2 Supervised Baseline
In addition to the unsupervised approach, we fine-tune XLM-R (base) Conneau et al. (2020) on the translation direction detection task to provide a supervised baseline for our experiments.
We extract the final hidden state for the first token by XLM-R for each segment of a translation, which serves as the feature representation for that segment. Inspired by the COMET architecture for MT evaluation Rei et al. (2020), we then combine the resulting representations for each translation by concatenating the addition of the representations of the segments of a pair, their absolute difference, and their product.555We deviate slightly from COMET by using the addition of the two segment representations instead of concatenating them sequentially, to avoid introducing input order bias.,666Furthermore, we experiment with encoding the source segment jointly with the translation. This yields similar results, but introduces input order bias. The test set results for this alternative architecture are listed in Appendix D. The resulting concatenation is used as input to train the classification head.
For the training set, we take a sample of 1398–1400 source segments and their corresponding HT and NMT-based translations per direction from WMT16. Additionally, we sample another 100 source segments per direction and their corresponding HT, NMT and pre-NMT-based translations from WMT16 as a validation set. We then train bilingual classifiers for each encs, enru and ende. We validate settings with different learning rates and epochs (Appendix C) and report the results for the test set in Table 2 with the setting that achieves the highest accuracies for the corresponding language pair777encs/ru: learning rate 1e-05, epoch 2; ende: learning rate 1e-05, epoch 4. on the validation set. The test set consists of HT and NMT sentence pairs from WMT21, WMT22 and WMT23, which is equivalent to the unsupervised experiments apart from the WMT16.
Since training the supervised systems on larger datasets could improve their performance, we experiment with training them on subsets of the Europarl corpus Koehn (2005); Ustaszewski (2019). Although higher accuracies are reached when the systems are tested on in-domain data, their performance drops below the WMT-based system when tested on out-of-domain data (WMT). Hence, we choose to describe the WMT-based model, which performs best on WMT data in the main part of this paper (Subsection 5.1) and give a detailed description and results of the Europarl-based system in Appendix E.
@@ 5.1 Sentence-level Classification
Language Pair | Avg. | ||
enru | 75.75 | 68.08 | 71.91 |
enuk | 72.76 | 72.56 | 72.66 |
enzh | 58.82 | 90.54 | 74.68 |
Macro-Avg. | 69.11 | 77.06 | 73.08 |
Table 3 compares the results for HT across all models. As a general result, we find that it is not NLLB, but M2M-100 that on average yields the best results for HT on the sentence level, with SMaLL-100 a close second. Hence, we report results for experiments with M2M-100 in the following, while performance of the other models is reported in Appendix A and B.
A second result is that the translation detection works best for NMT-based translations (75.0% macro-average), second-best for HT (66.5% macro-average), and worst for pre-NMT (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that the NMT systems we use tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems. In practice, one could pair our detection method with a monolingual model to identify such low-quality outputs.
In comparison to the results of the supervised system, we show in Table 2 that while our unsupervised approach is outperformed on HT, it is competitive for NMT-based translation. Taking into consideration the (cross-domain) results of the Europarl-based models shown in Table 17 as well, the main benefits of the unsupervised approach are highlighted: independence from training data and flexibility across languages and domains.
@@
5.2 LLM-generated Translations
The main focus of this paper lies on three different translation types: HT, NMT, and pre-NMT. However, LLMs have also shown strong translation capabilities (Kocmi et al., 2023). Hence, GPT-4 was considered as one of the translation systems in the WMT23 shared task and its outputs are therefore part of the NMT test set in this work. Table 6 shows the results of our translation direction detection approach on the GPT-4-generated test subset in isolation to document its performance on this additional translation type. For all three language pairs, our approach reaches accuracies that are comparable to the ones reached for the same pairs’ NMT translations.
5.3 Indirect Translations
With an experiment on the English-original FLORES data, we evaluate our approach on the special case that neither side is the original. As shown in Table 7, our approach yields relatively balanced predictions on human translations for csuk and xhzu, predicting each direction a roughly equal number of times. For defr, we again find that the model predicts defr more frequently than the reverse direction. The result for frzh display an even larger degree of disparity.
@@ 5.4 Directional Bias
When analyzing Table 3 for directional bias, we observe that M2M-100 is especially biased in the directions defr () and zhen (). While we expected a general bias towards xen due to the dominance of English in training data, we find that the direction and strength of the bias vary across language pairs and models. An extreme result is NLLB for enzh, with towards zhen.
There appears to be an inclination for bias towards languages closely related to English for some language pairs, such as defr and zhfr (Table 7) in contrast to language pairs, with both languages being closely related to (and including) English, such as enfr. The discrepancy in balance between dehi and zhfr indicates, however, that relatedness to English might not have a strong general influence on our approach, but rather that there is a bias for French and Chinese when applying it with M2M-100 or SMaLL-100.
Furthermore, the supervised systems exhibit higher bias (Table 2) than the unsupervised approach for all language pairs and translation types, indicating that directional bias is an issue for both supervised and unsupervised approaches.
We leave it to future work to explore whether bias can be reduced via different normalization strategies, a language pair specific bias correction term, or different model training. At present, our recommendation is to be mindful in the choice of NMT model and to perform validation before trusting the results of a previously untested NMT model for translation direction detection.
@@ 5.5 Qualitative Analysis
A qualitative comparison of sources and translations, as illustrated in Table 8, reveals that factors such as normalization, simplification, word order interference, and sentence length influence the detection of translation direction. In Example 1, an English HT translates the German source fully, while the NMT omits half of the content, showing a high degree of simplification. Our method recognizes the simplified NMT version as a translation, but not the more complete HT. Example 2 demonstrates normalization to a high degree. This shows that MMT-models assign extremely low probabilities when the target is more ungrammatical than the source. The third example indicates that translations exhibiting normalization, simplification, and interference to a higher degree are more likely to be identified. In Example 4, source language interference in terms of word order and choice significantly impacts the detection; the more literal translation mirroring the source’s word order is recognized, while the more liberal translation is not. Finally, Example 5 highlights challenges with short sentences: The German phrase Mit freundlichen Grüßen is fairly standardized, while its French equivalents can vary in use and context, adding to the ambiguity and affecting the probability distribution in NMT. Hence, our approach fails to identify any of the French translations without additional context.
@@
Furthermore, Table 8 shows that our approach more easily identifies translations that are simpler in terms of verbosity and sentence complexity, e.g.: Examples 2, 3, 4. While previous research indicates that verbosity is not the most reliable feature for supervised approaches Sominsky and Wintner (2019), it is likely a helpful feature for the unsupervised approach.
Misclassified short examples as in Example 5 are not a rarity in our experiments. Our findings show that an average accuracy comparable to that reported in Table 3 is attained starting at a sentence length between 60 and 70 characters.888We used SMaLL-100 for this analysis. See Appendix H. Additionally, we observed a trend where the accuracy of direction detection increases as the length of the sentences/documents grows. This aligns with previous unsupervised approaches, which also documented higher accuracy the larger the text chunks that were used, although there, reliable results were reported on a more extreme scale starting from text chunks with a length of 250 tokens Rabinovich and Wintner (2015).
5.6 Document-Level Classification
Accuracy scores for document-level results by M2M-100 (best-performing system at sentence level) are presented in Tables 9 (HT) and 10 (NMT). We consider documents with at least 10 sentences, and language pairs with at least 100 such documents in both directions.